OpenStudy (natasha18):

PLEASE HELP! Circle A has a radius five times longer than the radius of circle B. Which gives the ratio between the area of circle A and the area of circle B? 5 : 1 10 : 1 25 : 1 10π : 1 25π : 1

OpenStudy (peachpi):

\[\frac{ Area~A }{ Area~B }=\frac{ \pi (5r)^2 }{ \pi r^2 }=?\]